# presep= ]e1,e2,e3

### v1

@@ Appendix E Background on fundamental limit theorems
We introduce $\alpha$- and $\beta$-mixing using Rosenblatt (1956); Volkonskii and Rozanov (1959) as primary references (see also Bradley (2005); Kulik and Soulier (2020)) and state useful results on $\beta$-mixing sequences.
@@
Let $l\leq r$ and $m$ be integers. Let $\left(\Omega,\mathcal{F},\mathrm{Pr}\right)$ be a probability space on which the following sequences are defined, (i) a stationary $\mathbb{R}^{d}$-valued sequence $\left\{X_{j},j\in\mathbb{Z}\right\}$ with $\beta$-mixing coefficients $\beta_{j}$, (ii) a sequence $\left\{X_{j},j\in\mathbb{Z}\right\}$ such that the vectors $\left(X^{{\dagger}}_{(i-1)r+1},\ldots,X^{{\dagger}}_{ir}\right),i\in\mathbb{Z}$ are mutually independent and each have the same distribution as $\left(X_{1},\ldots,X_{r}\right)$, (iii) a sequence $\left\{\xi_{j},j\in\mathbb{Z}\right\}$ of i.i.d. random elements taking value in a measurable set $\left(E,\mathcal{E}\right)$, independent of the previous sequences. Let $Q_{m}$ and $Q_{m}^{{\dagger}}$ be the respective laws of $\left\{\left(\xi_{i},X_{(i-1)r+1},\ldots,X_{ir-l}\right),1\leq i\leq m\right\}$ and $\left\{\left(\xi_{i},X^{{\dagger}}_{(i-1)r+1},\ldots,X^{{\dagger}}_{ir-l}\right),1\leq i\leq m\right\}$. Then
@@
In this section we present existing results on regarding Central Limit Theorems using Billingsley (2009) and van der Vaart and Wellner (2023) as primary reference, see also Kulik and Soulier (2020).
@@
As the condition (20) is difficult to show directly, one often invokes the covering Central Limit Theorem when dealing with arrays of stochastic processes. In the theorem we below we present an adaption to the covering CLT in the case of separable processes.
@@
Assume $\mathbb{X}$ is a separable stochastic process indexed by the set $\mathcal{F}$. Then $||\mathbb{X}||_{\mathcal{F}}=\sup_{f\in\mathcal{F}}\{\mathbb{X}(f)\}$ is measurable.
@@
Denote by $\mathcal{P}$ the set of finite discrete probability measures on $E$. For $P\in\mathcal{P}$ define $L^{2}\left(P\right)=\left\{f:E\rightarrow\mathbb{R},P\left(f^{2}\right)<\infty\right\}.$ For each $f,g\in L^{2}\left(P\right)$ write
@@
The entropy condition (21) in Theorem 19 can be replaced with a uniform metric entropy condition with respect to the deterministic semi-metric $\mathrm{d}_{L^{2}\left(P\right)}$.
@@
Then condition (21) holds.
@@
If $\mathcal{G}$ is a VC-subgraph class with measurable envelope function $\boldsymbol{\mathrm{G}}$, then condition (23) holds. In particular, (23) holds for linearly ordered classes.
@@
[…]

### v2

@@ Appendix E Background on fundamental limit theorems
We introduce $\alpha$- and $\beta$-mixing using Rosenblatt1956ACL; Volkonskii1959 as primary references (see also Bradley_05; Kulik2020) and state useful results on $\beta$-mixing sequences.
@@
Let $l\leq r$ and $m$ be integers. Let $\left(\Omega,\mathcal{F},\mathrm{Pr}\right)$ be a probability space on which the following sequences are defined, (i) a stationary $\mathbb{R}^{d}$-valued sequence $\{X_{j},j\in\mathbb{Z}\}$ with $\beta$-mixing coefficients $\beta_{j}$, (ii) a sequence $\{X_{j},j\in\mathbb{Z}\}$ such that the vectors $(X^{{\dagger}}_{(i-1)r+1},\ldots,X^{{\dagger}}_{ir}),i\in\mathbb{Z}$ are mutually independent and each have the same distribution as $(X_{1},\ldots,X_{r})$, (iii) a sequence $\{\xi_{j},j\in\mathbb{Z}\}$ of i.i.d. random elements taking value in a measurable set $(E,\mathcal{E})$, independent of the previous sequences. Let $Q_{m}$ and $Q_{m}^{{\dagger}}$ be the respective laws of $\{(\xi_{i},X_{(i-1)r+1},\ldots,X_{ir-l}),1\leq i\leq m\}$ and $\{(\xi_{i},X^{{\dagger}}_{(i-1)r+1},\ldots,X^{{\dagger}}_{ir-l}),1\leq i\leq m\}$. Then
@@
In this section we present existing results on regarding Central Limit Theorems using BillingsleyPatrick2009Copm and vanderVaart2023 as primary reference, see also Kulik2020.
@@
As the condition of equation (17) is difficult to show directly, one often invokes the covering Central Limit Theorem when dealing with arrays of stochastic processes. In the theorem we below we present an adaption to the covering CLT in the case of separable processes.
@@
Assume $\mathbb{X}$ is a separable stochastic process indexed by the set $\mathcal{F}$. Then $\mathbb{X}||_{\mathcal{F}}=\sup_{f\in\mathcal{F}}\{\mathbb{X}(f)\}$ is measurable.
@@
Denote by $\mathcal{P}$ the set of finite discrete probability measures on $E$. For $P\in\mathcal{P}$ define $L^{2}\left(P\right)=\{f:E\rightarrow\mathbb{R},P(f^{2})<\infty\}.$ For each $f,g\in L^{2}\left(P\right)$ write
@@
The entropy condition of equation (18) in Theorem 19 can be replaced with a uniform metric entropy condition with respect to the deterministic semi-metric $\mathrm{d}_{L^{2}\left(P\right)}$.
@@
Then equation (18) holds.
@@
If $\mathcal{G}$ is a VC-subgraph class with measurable envelope function $\boldsymbol{\mathrm{G}}$, then (20) holds. In particular, (20) holds for linearly ordered classes.
@@
Consider non-empty intervals $I,J$ and let $\left\{\mathbb{X},\mathbb{X}_{n},n\geq 1\right\}$ be $\mathbb{D}\left(I\right)$-valued stochastic processes and assume that $\mathbb{X}$ has almost surely continuous sample paths. Let $h:I\rightarrow J$ be a surjective, continuous and increasing function with positive derivative $h^{\prime}$. If there exists a deterministic sequence $c_{n}\to\infty$ such that $c_{n}\left(\mathbb{X}_{n}-h\right)\overset{d}{\to}\mathbb{X}$ in $\mathbb{D}\left(I\right)$ endowed with the $J_{1}$-topology and $\mathbb{X}_{n},n\geq 1$ have non-decreasing sample paths, then
$c_{n}\left(\mathbb{X}_{n}-h,\mathbb{X}^{\leftarrow}_{n}-h^{\leftarrow}\right)\overset{d}{\to}\left(\mathbb{X},-\left(h^{\leftarrow}\right)^{\prime}\times\mathbb{X}\circ h^{\leftarrow}\right),$